For the standard normal distribution, find the area between -1.55 and 1.90.


1
Expert's answer
2021-03-01T15:22:14-0500

"P(-1.55<z<1.90)\\
=P(0<z<-1.55)+P(0<z<1.90)\\
=0.4394+0.4713=0.9107"